24.36 mL of .1000M NaOH is reacted with 50.00 mL of .1000M HCl reacted to produce a NaCl solution. If the density of the resulting solution is 1.224 g/mL, what is the mass of the NaCl solution produced?
Answer #1

Total volume of solution = 24.36 + 50 = 74.36 ml

Density of NaCl solution = 1.224 g/mL

So mass of NaCl solution = density * volume

= 74.36 * 1.224

= 91 g